OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (3, 7).

OpenStudy (anonymous):

\[y-7=2(x-3)\]

OpenStudy (amistre64):

\[y = 2(x-3)+7\] \[y = 2x -6+7\] \[y = 2x +1\]